karen and andrew start out from towns 164 km apart and drive towards each other. Andrew travels at a rate of 40km/h and karen at 48km/h. If karen begins the journey 15 minutes after andrew, how long after andrew leaves will they meet?
Answer by lwsshak3(11628)

karen and andrew start out from towns 164 km apart and drive towards each other. Andrew travels at a rate of 40km/h and karen at 48km/h. If karen begins the journey 15 minutes after andrew, how long after andrew leaves will they meet?
let x=travel time of karen
15 min=1/4 hr
x+1/4 =travel time ofAndrew
Dstance=travel time*rate
48x+40(x+1/4)=164
48x+40x+10=154
88x=154
x=31.75hr=1 hr and 45 min
how long after andrew leaves will they meet? 1 hr and 45 min